# The local moduli of Sasaki-Einstein $\mathbb{Q}$-homology 7-spheres and invertible polynomials

### v1

@@ Abstract.
We compute the dimension of the local moduli space of Sasaki-Einstein metrics for links of invertible polynomials coming from the list of Johnson and Kollár of anticanonically embedded Fano 3-folds of index 1 that produce $\mathbb{Q}$-homology 7-spheres, that is, 7-manifolds whose $\mathbb{Q}$-homology equals that of the 7-sphere. In order to do so, we propose additional conditions to the Diophantine equations associated to this problem. We also find solutions for the problem associated to the moduli for the Berglund-Hübsch duals of links arising from Thom-Sebastiani sums of chain and cycle polynomials. In particular we give an explicit description of local moduli of the complex structure for the corresponding Kähler-Einstein orbifold which in this setting is a ${\mathbb{Q}}$-homology projective 3-spaces with quotient singularities.
Keywords: Moduli, Sasaki-Einstein metrics, Rational homology spheres, Berglund-Hübsch.
@@ 1. Introduction
The moduli problem, that is, the study of spaces that classify geometric structures on manifolds up to isomorphism, is highly relevant in Geometry. In particular the moduli of Einstein structures has received much attention, see Chapter 12 of [2]. In many cases, finding a space with suitable properties that parametrizes the isomorphism classes is too hard and even is the space is theoretically described, sometimes in practice it is not possible to provide effective tools to compute the moduli space on explicit examples. However, the local problem sometimes is easier to handle. This was the approach taken by Kodaira and Spencer in their canonical work on deformations of complex structures on compact manifolds [25].
@@
For Sasakian manifolds, which are roughly speaking the odd dimensional analogue of Kähler manifolds, the moduli problem has been addressed and an appropriate notion of moduli space has been achieved, see [3] and references therein (c.f.[31]). In particular, finding the number of components of this moduli has been studied and used to obtain lower bounds for the dimension of the moduli space for links of Brieskorn-Pham polynomials and Smale manifolds [10, 20, 7, 26]. For instance, the moduli space of Sasakian metrics with positive Ricci curvature on odd homotopy spheres bounding parallelizable manifolds has infinitely many positive components [8]. An important ingredient to obtain information on the moduli is given by the use of invertible polynomials, particularly to describe the moduli of Sasaki-Einstein metrics.
@@
Indeed, invertible polynomials have been studied in the context of Sasakian geometry for some time. These polynomials, thought as defining subsets of weighted projective spaces, produce links with some interesting features both at topological/differential and Riemannian level. As a matter of fact, in [4] Boyer and Galicki generalized a method introduced by Kobayashi in [22] to produce Sasaki-Einstein metrics on links of hypersurface singularities. From there, many examples of highly connected Sasaki-Einstein manifolds were found. For such examples the existence of orbifold Kähler-Einstein metrics on various Fano orbifolds is verified using different incarnations of the $\alpha$-invariant method developed by Tian, Nadel, Demailly and Kollár [35, 29, 16]. In [20, 6, 7, 11] special interest was given to 5-dimensional manifolds and homotopy spheres realized as links of Brieskorn-Pham polynomials which are an important block in the classification of invertible polynomials. In these articles the local moduli space of Sasaki-Einstein structures for these links were also studied. For instance in [7] they exhibited homotopy spheres given as links of dimension $32n-$ 3 such that the number of inequivalent families of Sasaki-Einstein metrics and the dimension of their moduli grows double exponentially with the dimension and they also showed that all the 28 oriented diffeomorphism classes on $S^{7}$ admit inequivalent families of Sasakian-Einstein structures. Later, in [12] it is shown that the moduli space of positive Sasaki classes with vanishing first Chern class for manifolds of the form $S^{2n}\times S^{2n+1}$ or for odd dimensional homotopy spheres among others, has a countably infinite number of components of dimension greater than one and these contain no extremal Sasaki metrics at all. Again this information is produced through links of Brieskorn-Pham polynomials.
@@
In [5], based on the complete list of Johnson and Kollár of anticanonically embedded Fano 3-folds of index one [19], many $\mathbb{Q}$-homology 7-spheres admitting Sasaki-Einstein metrics were constructed using the Kobayashi-Boyer-Galicki method. In [13] the authors were able to enlarge the previous list of examples due to a result [18] on the validity of an Orlik’s conjecture [32] on how to compute the homology of links of invertible polynomials. In [14], in joint work with R. Gomez, we studied these examples in a more systematic way and showed that this list of $\mathbb{Q}$-homology 7-spheres remain $\mathbb{Q}$-homology 7-spheres under the Berglund-Hübsch transpose rule from classical mirror symmetry [1] and under very mild conditions all of them leave the Einstein condition invariant. We found that all the invertible polynomials producing $\mathbb{Q}$-homology 7-spheres and admitting Sasaki-Einstein metrics were polynomials of cycle type, or iterated Thom-Sebastiani sums of cycle and chain types and not of Brieskorn-Pham type. It is natural to determine the family of inequivalent Sasaki-Einstein on each of these $\mathbb{Q}$-homology 7-spheres. In this article we study the space of deformations of Sasaki-Einstein metrics for this sort of links. We found that only the latter case admits non-equivalent Sasaki-Einstein metrics. Actually we show that
@@
  • For polynomials of cycle type producing $\mathbb{Q}$-homology 7-spheres the corresponding local moduli space of Sasaki-Einstein metrics has dimension zero, that is, links of this sort do not admit inequivalent families of Sasaki-Einstein structures, see Subsection 3.1.
  • For Thom-Sebastiani sums of chain and cycle polynomials the generators of local moduli space of Sasaki-Einstein metrics depend merely on a few parameters that are given in terms of the exponents of the polynomial and the decomposition of the corresponding weights, see Subsection 3.2.
  • For the Berglund-Hübsch dual link, where one only needs to study polynomials of chain-cycle type of Fano index greater than one (see Subsection 3.3), we found that $\mathbb{Q}$-homology 7-spheres given as links coming from polynomials of this type do not admit inequivalent families of Sasaki-Einstein structures with the exception of five elements.
@@
In order to achieve these results, we find the generators of the infinitesimal deformations of the complex structures for orbifolds arising from families of polynomials of the types mentioned above (see Subsection 2.2). In particular our results give information on the connected components of the moduli of $\mathbb{Q}$-homology 7-spheres with specific differential structure which can be expressed as link of isolated hypersurface singularities coming from Johnson and Kollár’s list of Fano 3-folds of index 1. Our findings can be interpreted in two different setting:
@@
  • Seifert $S^{1}$-bundles are $\mathbb{Q}$-homology spheres if and only if the corresponding orbifolds are $\mathbb{Q}$-homology complex projective spaces [21], so our results describes some components of the moduli space of ${\mathbb{Q}}$-homology complex projective 3-spaces with quotient singularities.
@@
  • Sasaki-Einstein structures on the manifold determine Ricci-flat Kähler cone metrics on the corresponding affine cone [8], so our results give information on the moduli of Calabi-Yau cones [31].
The general strategy is to find the arithmetic conditions provided by the invertible polynomials $f$ that produce links $L_{f}$ that are $\mathbb{Q}$-homology 7 spheres and also admit Sasaki-Einstein metrics. Then we interpret these arithmetic conditions as constraint to obtain all possible monomials
@@
of certain degree, that describe the space of deformations of holomorphic structures on the orbifold $Z_{f}=\{f=0\}/\mathbb{C}^{*}$ associated to the link. Thus in order to find the number of monomials we obtain a Diophantine equation subject to certain the conditions that we are able to solve for the different invertible polynomials that appear in this problem.
The paper is organized as follows. Section 2 reviews the background material relevant for this paper. In Section 3 we relate the problem of finding the moduli number of Sasaki-Einstein $\mathbb{Q}$-homology 7-spheres to solving the Diophantine equations subject to certain arithmetic conditions extracted from the building blocks that conform the invertible polynomial associated to this problem; then we find the dimension of the corresponding local moduli space of Sasaki-Einstein metrics and prove Theorems 3.1.1, 3.2.1, 3.3.1.
@@ 2. Preliminaries
2.1. Sasakian structures on links and the Berglund-Hübsch rule
Below we give the basics on Sasakian geometry and list some important results that we will need, the canonical references here are [8] and [34]. Then we focus on links of invertible polynomials with certain arithmetic conditions which produce $\mathbb{Q}$-homology 7-spheres and explain how the Berglund-Hübsch transpose rule can be used in this framework.
Sasakian Geometry: Sasakian geometry is a special type of contact metric structure on a $(2n+1)$-dimensional manifold $M$ described by the quadruple of tensors $(\Phi,\xi,\eta,g)$ such that $\eta$ is a contact 1-form, $\Phi$ is an endomorphism of the tangent bundle, $g=d\eta\circ(\mathbb{I}\times\Phi)+\eta\otimes\eta$ is a Riemannian metric and $\xi$ is the Reeb vector field which is Killing. Moreover, the underlying CR-structure $\left(\mathcal{D},\left.\Phi\right|_{\mathcal{D}}\right)$ is integrable, where $\mathcal{D}=\ker\eta.$
A Sasaki manifold has a transverse Kählerian structure $(\nu(\mathcal{F}_{\xi}),\bar{J})$ determined by the normal bundle $\nu(\mathcal{F}_{\xi})$ of the characteristic foliation $\mathcal{F}_{\xi}$ determined by $\xi$ and the natural complex structure $\bar{J}$ in $\nu(\mathcal{F}_{\xi})$ induced by $\left.\Phi\right|_{\mathcal{D}}.$ In fact, when all the orbits of $\xi$ are closed, the Reeb vector field $\xi$ generates a locally free circle action whose quotient is a Kähler orbifold. In this case the Sasakian structure is called quasi-regular. When the action is free it is called regular and the quotient is a Kähler manifold. In the irregular case when the orbits of the Reeb vector field are not closed the local quotients are Kähler.
On the other hand, if we consider the cone $C(M)=M\times\mathbb{R}^{+}$ with symplectic form $\omega=$ $d\left(r^{2}\eta\right)$ where $r$ is the radial coordinate. Using the Liouville vector field $\Psi=r\partial_{r}$ we define a natural complex structure $I$ on $C(M)$ by
@@ 2.1. Sasakian structures on links and the Berglund-Hübsch rule
where $X$ is a vector field on $M$, and $\xi$ is understood to be lifted to $C(M)$. By adding the apex of the cone we obtain an affine variety $Y=C(M)\cup\{0\}$ which has been intensely studied to obtain K-stability theorems in the manner of the work of Chen-Donaldson-Sun in the context of Sasakian manifolds [15]. We know that $(M,g)$ is Sasakian if the cone $(C(M),\omega,I,\bar{g})$ is Kähler with Kähler metric given by $\bar{g}=dr^{2}+r^{2}g.$ Moreover, the Sasakian structure corresponds to a polarized affine variety $(Y,I,\xi)$ polarized by the Reeb vector field $\xi$ and this affine variety admits a Ricci-flat Kähler cone metric if and only if $(M,g)$ admits a Sasaki-Einstein metric.
Let us denote by $\mathcal{S}(M)$ the space of all Sasakian structures on $M$ and give it the $C^{\infty}$ Fréchet topology as sections of vector bundles, and denote by $\mathcal{S}(M,\xi,\bar{J})$ the subset of Sasakian structures with Reeb vector field $R$ and transverse holomorphic structure $\bar{J}$. The set $\mathcal{S}(M,\xi,\bar{J})$ is endowed with the subspace topology. Consider $\mathcal{SE}(M)$ the subspace of Sasaki-Einstein metrics in $\mathcal{S}(M,\xi,\bar{J})$ and let $\mathfrak{Aut}(\bar{J})$ be the group of complex automorphisms of $(C(M),I)$ that commute with $\Psi-i\xi$. This group descends to an action on $(M,\mathcal{S})$ commuting with $\xi.$ Denote by $\mathfrak{Aut}(\bar{J})_{0}$ the connected component of $\mathfrak{Aut}(\bar{J}).$ Then one defines the local moduli space of Sasaki-Einstein metrics [3] by
$\mathfrak{M}^{SE}(M)=\mathcal{S}\mathcal{E}(M)/\mathfrak{Aut}(\bar{J})_{0}.$
As we will see in Subsection 2.2, this space has a very concrete description for links of invertible polynomials.
An important source of examples of Sasakian manifolds in a very general way can be given as total spaces of $S^{1}$-orbibundles over projective orbifolds. As a special case, one can build up Sasakian structures on links of hypersurface singularities of weighted homogeneous polynomials and we explain how to do this below.
Let us recall that a polynomial $f\in\mathbb{C}\left[z_{0},\ldots,z_{n}\right]$ is said to be a weighted homogeneous polynomial of degree $d$ and weight $\mathbf{w}=$ $\left(w_{0},\ldots,w_{n}\right)$, $w_{i}\in\mathbb{Z}^{+}$, and if for any $\lambda\in\mathbb{C}^{*}$
@@
is smooth everywhere except at the origin in $\mathbb{C}^{n+1}$ which is equivalent to assume that the weighted variety
$(V_{f}-\{0\})/\mathbb{C}^{*}\subset\mathbb{P}(\mathbf{w})$
is quasi-smooth of dimension $n.$
A link $L_{f}(\mathbf{w},d)$ of a hypersurface singularity is defined as $f^{-1}(0)\cap S^{2n+1}$, where $S^{2n+1}$ is the $(2n+1)$-sphere in Euclidean space. By the Milnor fibration theorem [27], $L_{f}(\mathbf{w},d)$ is a closed ($n-2)$-connected manifold that bounds a parallelizable manifold with the homotopy type of a bouquet of $n$-spheres. Furthermore, $L_{f}(\mathbf{w},d)$ admits a quasi-regular Sasaki structure $\mathcal{S}=\left(\xi_{\mathbf{w}},\eta_{\mathbf{w}},\Phi_{\mathbf{w}},g_{\mathbf{w}}\right)$ which is the restriction of the weighted Sasakian structure on the sphere $S^{2n+1}$ with Reeb vector field $\xi_{\mathrm{w}}=\sum_{k=0}^{n}w_{k}\left(y_{k}\partial_{{x}_{k}}-x_{k}\partial_{{y}_{k}}\right).$ Moreover, if one considers the locally free $S^{1}$-action induced by the weighted $\mathbb{C}^{*}$ action on $f^{-1}(0)$ the quotient space of the link $L_{f}(\mathbf{w},d)$ by this action is the weighted hypersurface $Z_{f},$ a Kähler orbifold: we have the following commutative diagram [8]

### v2

# The local moduli of Sasaki-Einstein rational homology 7-spheres and invertible polynomials

@@ Abstract.
We study the local moduli space of Sasaki-Einstein metrics on links of invertible polynomials defining rational homology 7-spheres. All these polynomials are either of cycle type or are given as Thom Sebastiani sums of a cycle block and another atomic block. We found that for polynomials of cycle type, the local moduli spaces of Sasaki-Einstein metrics are zero dimensional. For the Thom-Sebastiani sums of an atomic block and a cycle polynomial, the dimensions of the local moduli spaces of Sasaki-Einstein metrics are positive in general. Since all the links under study in this article remain Sasaki-Einstein rational homology 7-spheres under the Berglund-Hübsch rule from classical mirror symmetry [1, 16], we are able to find solutions for the problem associated to the moduli for the Berglund-Hübsch transpose duals of this type of links. For the purpose of doing this, we give specific description of the moduli spaces of complex structures on the weighted quasismooth hypersurfaces cut out by the corresponding invertible polynomials and, in particular, from this description, we can produce families of quasismooth weighted hypersurfaces that denegerate to non-quasismooth with at worst klt singularties.
Keywords: Moduli, Sasaki-Einstein metrics, Rational homology spheres, Berglund-Hübsch, Calabi-Yau cones
@@ 1. Introduction
For Sasakian manifolds, which are roughly speaking the odd dimensional analogue of Kähler manifolds, the moduli problem has been addressed and an appropriate notion of moduli space has been designed, however little is known about this space, see [3] and references therein. In particular, finding the number of components of this moduli has been studied in order to obtain lower bounds for the dimension of the moduli space of Sasakian structures for links of Brieskorn-Pham polynomials and Smale manifolds [11, 28, 7]. For Sasaki-Einstein structures, the local study on its moduli has given interesting results via deformations of the transverse complex structure in a more general setting, see [40, 48].
In this article, we describe the local moduli of Sasaki-Einstein metrics from the study of invertible polynomials, that is polynomials of the form $f=\sum_{i=1}^{n}\prod_{j=1}^{n}x_{j}^{a_{ij}},$ where $A=\left(a_{ij}\right)_{i,j=1}^{n}$ is a non-negative integer-valued matrix which is invertible over $\mathbb{Q}$ and where $f$ is quasihomogeneous, i.e., there exist positive integers $w_{j}$ such that $d:=\sum_{j=1}^{n}w_{j}a_{ij}$ is constant for all $i$, and $f$ is quasismooth, i.e., $f:\mathbb{C}\left[x_{1},\ldots,x_{n}\right]\rightarrow\mathbb{C}$ has exactly one critical point at the origin. Due to the Kreuzer-Skarke classification of invertible polynomials [32] we know that any invertible polynomial, up to permutation of variables, can be written as a Thom-Sebastiani sum of three types of polynomials usually called atoms:
  1. (1)
    Fermat type: $w=x^{a}$,
  2. (2)
  3. (3)
@@
These atoms have remarkable properties, for instance in [29] cyclic polynomials have been studied in order to obtain information on rational surfaces with quotient singularities. Invertible polynomials, thought as defining subsets of weighted projective spaces, produce smooth links with some interesting features both at the topological/differential and the Riemannian level. As a matter of fact, in [4] Boyer and Galicki generalized a method introduced by Kobayashi in [26] to produce Sasaki-Einstein metrics on links of hypersurface singularities. From there, many examples of highly connected Sasaki-Einstein manifolds were found. For such examples the existence of orbifold Kähler-Einstein metrics on various Fano orbifolds is verified using different incarnations of the $\alpha$-invariant method developed by Tian, Nadel, Demailly, Kollár and others [46, 39, 19, 18]. In [28, 6, 7, 12] special interest was given to 5-dimensional manifolds and homotopy spheres realized as links of Brieskorn-Pham polynomials which are an important block in the classification of invertible polynomials. In these articles the local moduli space of Sasaki-Einstein structures for this kind of links were also studied. For instance, in [7] it was found that for homotopy spheres given as links, the dimension of their moduli grew double exponentially with the dimension and they also showed that all the 28 oriented diffeomorphism classes on $S^{7}$ admitted inequivalent families of Sasakian-Einstein structures. In [8], again using links of Brieskorn-Pham polynomials viewed as branched covers over Calabi-Yau hypersurfaces, it is shown that there exist continuous parameter families of Sasakian-Einstein metrics on infinitely many simply connected rational homology spheres in every odd dimension greater than 3. In particular, for dimension 7 they found 38 effective real parameters. Later, in [13] it is shown that the moduli space of positive Sasaki classes with vanishing first Chern class for manifolds of the form $S^{2n}\times S^{2n+1}$ or for odd dimensional homotopy spheres among others, has a countably infinite number of components of dimension greater than one and these contain no extremal Sasaki metrics at all. Again this information is produced through links of Brieskorn-Pham polynomials.
@@
In [5], based on the complete list of Johnson and Kollár of anticanonically embedded Fano 3-folds [24], many rational homology 7-spheres admitting Sasaki-Einstein metrics were constructed using the Kobayashi-Boyer-Galicki method. The ideas in [5] were extended in [14] and [16] and moreover it was shown that all the invertible polynomials producing rational homology 7-spheres admitting Sasaki-Einstein metrics were polynomials of cycle type, or iterated Thom-Sebastiani sums of a cycle type involving three variables and another atomic type involving two variables and they cannot be written as Brieskorn-Pham polynomials. In this article we investigate the space of deformations of Sasaki-Einstein metrics for this sort of links. We found that only the latter case admits non-equivalent Sasaki-Einstein metrics. Actually we show that
@@
  • For polynomials of cycle type producing rational homology 7-spheres, the corresponding local moduli space of Sasaki-Einstein metrics has dimension zero, that is, links of this sort do not admit inequivalent families of Sasaki-Einstein structures, see Subsection 3.1.
  • For Thom-Sebastiani sums of a cycle type and another atomic type, the dimension of local moduli space of Sasaki-Einstein metrics are given in terms of the rational weights, introduced by Milnor and Orlik in [38], for the corresponding weighted hypersurface and, in general, this dimension is positive, see Subsection 3.2.
@@
Since all the links under study in this article remain Sasaki-Einstein rational homology 7-spheres under the Berglund-Hübsch rule from classical mirror symmetry [16], we found that rational homology 7-spheres given as links coming from polynomials of this type do not admit inequivalent families of Sasaki-Einstein structures with the exception of five elements, see Subsection 3.3. In particular our results give information on the connected components of the moduli of rational homology 7-spheres with specific differential structure which can be expressed as link of isolated hypersurface singularities coming from the Johnson and Kollár list of Fano 3-folds of index 1. Our findings can be interpreted in two different settings:
@@
  • Seifert $S^{1}$-bundles are rational homology spheres if and only if the corresponding orbifolds are rational homology complex projective spaces [30], so our results describe some components of the moduli space of rational homology complex projective 3-spaces with quotient singularities.
@@
  • Sasaki-Einstein structures on the manifold determine Ricci-flat Kähler cone metrics on the corresponding affine cone [9], so our results give information on the moduli of Calabi-Yau cones.
In order to obtain these results, we find the generators of the infinitesimal deformations of the complex structures for orbifolds arising from families of polynomials of the types mentioned above, and then determine the moduli problem for Kähler-Einstein metrics on the corresponding hypersurfaces embedded in weighted projective spaces. In fact, arithmetic conditions on the invertible polynomials are imposed so the corresponding links are rational homology 7-spheres admitting Sasaki-Einstein metrics. Then we we use these arithmetic constraints to obtain all possible monomials
@@
of certain degree $d.$
Due to our interest in smooth Sasaki-Einstein links, our results are based on the information obtained on the subset all quasismooth elements determined by the monomials in $H^{0}(\mathbb{P}(\mathbf{w}),\mathcal{O}(d)).$ Nonetheless, the explicit description of $H^{0}(\mathbb{P}(\mathbf{w}),\mathcal{O}(d))$ leads to the study of the non-quasismooth polynomial generated by these monomials where, as suggested in [34] (see also [41]), the ones with at worst klt singularities are the candidates to give rise to non-smooth links whose metric cones can be considered as some sort of degenerating Calabi-Yau cones. Actually, non-quasismooth klt hypersurfaces with $\alpha$-invariant greater than one describe points in the boundary of a compactification of the moduli of quasismooth Kähler-Einstein weighted hypersurface and on the $K$-moduli of Fano cones as described in [42] for instance. We explore these ideas in the last section of this article and explain how to produce families of quasismooth weighted hipersurfaces that denegerate to non-quasismooth klt varieties through examples.
The paper is organized as follows. Section 2 reviews the background material relevant for this paper. In Section 3 we relate the problem of finding the moduli number of Sasaki-Einstein rational homology 7-spheres to solving the Diophantine equations subject to certain arithmetic conditions extracted from the building blocks of the invertible polynomial associated to this problem. Finally, we find the dimension of the corresponding local moduli space of Sasaki-Einstein metrics and prove our main results. In the last section, we discuss the role of non-quasismooth hypersurfaces in this setting.
@@ 2. Preliminaries
2.1. Sasakian structures on smooth links and invertible polynomials
Below we give a very brief review of the main ingredients we need to establish our results. The canonical references here are [9] and [45]. After that, we give arithmetic conditions that allow us to manufacture rational homology 7-spheres from links of invertible polynomials. Finally, we explain how the Berglund-Hübsch transpose rule [1] can be used in this framework.
Sasakian Geometry: Sasakian geometry is a special type of contact metric structure on a $(2n+1)$-dimensional manifold $M$ described by the tensors $(\Phi,\xi,\eta,g)$ such that $\eta$ is a contact 1-form, $\Phi$ is an endomorphism of the tangent bundle, $g=d\eta\circ(\mathbb{I}\times\Phi)+\eta\otimes\eta$ is a Riemannian metric and $\xi$ is the Reeb vector field which is Killing. Moreover, the underlying CR-structure $\left(\mathcal{D},\left.\Phi\right|_{\mathcal{D}}\right)$ is integrable, where $\mathcal{D}=\ker\eta$ denotes the contact structure.
A Sasaki manifold has a transverse Kählerian structure $(\nu(\mathcal{F}_{\xi}),\bar{J})$ determined by the normal bundle $\nu(\mathcal{F}_{\xi})$ of the characteristic foliation $\mathcal{F}_{\xi}$ determined by $\xi$ and the natural complex structure $\bar{J}$ in $\nu(\mathcal{F}_{\xi})$ induced by $\left.\Phi\right|_{\mathcal{D}}.$ In fact, when all the orbits of $\xi$ are closed, the Reeb vector field $\xi$ generates a locally free circle action whose quotient is a Kähler orbifold. In this case the Sasakian structure is called quasiregular. When the action is free it is called regular and the quotient is a Kähler manifold. In the irregular case when the orbits of the Reeb vector field are not closed the local quotients are Kähler.
Alternatively, one can understand a Sasakian structure on the manifold $M$ in terms of the metric cone $C(M)=M\times\mathbb{R}^{+}$ with symplectic form $\omega=$ $d\left(r^{2}\eta\right)$ where $r$ is the radial coordinate. Indeed, using the Liouville vector field $\Psi=r\partial_{r}$ we define a natural complex structure $I$ on $C(M)$ by
@@ 2.1. Sasakian structures on smooth links and invertible polynomials
where $X$ is a vector field on $M$, and $\xi$ is understood to be lifted to $C(M)$. By adding the apex of the cone we obtain an affine variety $C(M)\cup\{0\}$ which has been intensely studied to obtain K-stability theorems in the manner of the work of Chen-Donaldson-Sun in the context of Sasakian manifolds [17]. We know that $(M,g)$ is Sasakian if the cone $(C(M),\omega,I,\bar{g})$ is Kähler with Kähler metric given by $\bar{g}=dr^{2}+r^{2}g.$ Moreover, the Sasakian structure corresponds to a polarized affine variety $(C(M)\cup\{0\},I,\xi)$ polarized by the Reeb vector field $\xi$ and this affine variety admits a Ricci-flat Kähler cone metric if and only if $(M,g)$ admits a Sasaki-Einstein metric.
Let us denote by $\mathcal{S}(M)$ the space of all Sasakian structures on $M$ with the $C^{\infty}$ Fréchet topology as sections of vector bundles, and let $\mathcal{S}(M,\xi,\bar{J})$ be the subset of Sasakian structures with Reeb vector field $\xi$ and transverse holomorphic structure $\bar{J},$ which is endowed with the subspace topology. Consider $\mathcal{SE}(M)$ the subspace of Sasaki-Einstein metrics in $\mathcal{S}(M,\xi,\bar{J})$ and let $\mathfrak{Aut}(\bar{J})$ be the group of complex automorphisms of $(C(M),I)$ that commute with $\Psi-i\xi$. This group descends to an action on $(M,\mathcal{S})$ commuting with $\xi.$ Then one defines the local moduli space of Sasaki-Einstein metrics [3] by
$\mathfrak{M}^{SE}(M)=\mathcal{S}\mathcal{E}(M)/\mathfrak{Aut}(\bar{J})_{0},$
where $\mathfrak{Aut}(\bar{J})_{0}$ denotes the connected component of $\mathfrak{Aut}(\bar{J}).$
As we will see in Subsection 2.2, this space has a very concrete description for links of weighted homogeneous hypersurfaces.
Weighted homogeneous hypersurfaces: Consider the weighted $\mathbb{C}^{*}$-action on the affine space $\mathbb{C}^{n+1},$ defined by
$\left(z_{0},\ldots,z_{n}\right)\mapsto\left(\lambda^{w_{0}}z_{0},\ldots,\lambda^{w_{n}}z_{n}\right)$
where $\mathbf{w}=\left(w_{0},\ldots,w_{n}\right)$ is a sequence of positive integers. Then we obtain the weighted projective space with a canonical orbifold structure, defined as the quotient space
$\mathbb{P}(\mathbf{w})=\left(\mathbb{C}^{n+1}-\{\mathbf{0}\}\right)/\mathbb{C}^{*}.$
The weighted projective space $\mathbb{P}(\mathbf{w})$ is said to be well-formed if the weighted $\mathbb{C}^{*}$-action on $\mathbb{C}^{n+1}$ has trivial stabilizers in codimension 1, that is, when $\operatorname{gcd}\left(w_{0},\ldots,\widehat{w}_{i},\ldots,w_{n}\right)=1$ for each $i.$ Here the hat symbol means delete that corresponding element. As an algebraic variety, a weighted projective space can be defined as
$\mathbb{P}\left(w_{0},\ldots,w_{n}\right)=\operatorname{Proj}(S(\mathbf{w})),$
where $S(\mathbf{w})=\mathbb{C}\left[x_{0},\ldots,x_{n}\right]$ is the graded polynomial ring such that the weight of each $x_{i}$ equals $w_{i}.$
Let us recall that a polynomial $f\in\mathbb{C}\left[z_{0},\ldots,z_{n}\right]$ is said to be a weighted homogeneous polynomial of degree $d$ and weight vector $\mathbf{w}=$ $\left(w_{0},\ldots,w_{n}\right)$, if for any $\lambda\in\mathbb{C}^{*}$
@@
is smooth everywhere except at the origin in $\mathbb{C}^{n+1}$ which is equivalent to say that the weighted variety
$Z_{f}=(V_{f}-\{\mathbf{0}\})/\mathbb{C}^{*}\subset\mathbb{P}(\mathbf{w})$
is quasismooth of dimension $n.$ There are well-known conditions that determine when a specific polynomial determines a quasismooth weighted hypersurface [21, 24]:
Lemma 2.1.1.
A weighted hypersurface of degree $d$ in $\mathbb{P}\left(w_{0},\ldots,w_{4}\right)$, where $d>w_{i}$, is quasismooth if and only if the following hold:
  1. (1)
    For each $i=0,\cdots,4$ there is a $j$ and a monomial $z_{i}^{m_{i}}z_{j}$ of degree $d$. Here $j=i$ is possible.
  2. (2)
    For all distinct $i,j$ either there is a monomial $z_{i}^{b_{i}}z_{j}^{b_{j}}$ of degree $d$ or there exist monomials $z_{i}^{n_{1}}z_{j}^{m_{1}}z_{k},z_{i}^{n_{2}}z_{j}^{m_{2}}z_{l}$ of degree $d$ with $\{k,l\}\neq\{i,j\}$ and $k\neq l$.
  3. (3)
    For every $i,j$ there exists a monomial of degree $d$ that does not involve either $z_{i}$ or $z_{j}$.
We have the following important result whose proof can be found in [9], Corollary 5.4.8. First, recall that the index $I$ of a weighted hypersurface is given by the difference $I=|\mathbf{w}|-d,$ where $|\mathbf{w}|=\sum_{i=0}^{n}w_{i}$ denotes the norm of the weight vector $\mathbf{w}.$ The weighted hypersurface is said to be Fano if $I>0.$
Theorem 2.1.1.
Let $Z_{f}\subset\mathbb{P}(w_{0},\ldots w_{n})$ be a quasismooth weighted homogeneous Fano hypersurface of degree $d$. Then $Z_{f}$ admits a Kähler-Einstein orbifold metric if the following estimate holds:
$dI<\frac{n}{(n-1)}\min_{i,j}\left\{w_{i}w_{j}\right\}.$
The weighted hypersurface is well-formed if $\mathbb{P}(\mathbf{w})$ is well-formed and $Z_{f}\cap{\text{sing}}(\mathbb{P}(\mathbf{w}))$ has codimension at least 2 in $Z_{f}$. When $Z_{f}$ is well-formed, the canonical divisor satisfies the adjunction formula $K_{Z_{f}}=\mathcal{O}_{Z_{f}}\left(d-w_{0}-\cdots-w_{n}\right)$. In [21] a criterion to determine the well-formedness of the weighted variety is given: a hypersurface defined by the weighted homogeneous polynomial $f$ of degree $d$ is well-formed in the well-formed weighted projective space $\mathbb{P}\left(w_{0},\ldots,w_{n}\right)$ if $\operatorname{gcd}\left(w_{0},\ldots,\hat{w}_{i},\ldots,\hat{w}_{j},\ldots,w_{n}\right)\mid d$ for distinct $i,j=0,\ldots,n$.
Links of hypersurface singularities and Sasakian structures. Sasakian structures can be manufactured on links of hypersurface singularities of weighted homogeneous polynomials and we explain how to do this below.
A link $L_{f}(\mathbf{w},d),$ or $L_{f}$ for short, is defined as the intersection $V_{f}\cap S^{2n+1}$, where $S^{2n+1}$ is the $(2n+1)$-sphere in Euclidean space. By the Milnor fibration Theorem [37], $L_{f}(\mathbf{w},d)$ is a closed (n2)n-2)-connected manifold that bounds a parallelizable manifold with the homotopy type of a bouquet of $n$-spheres. Furthermore, $L_{f}(\mathbf{w},d)$ admits a quasiregular Sasaki structure $\mathcal{S}=\left(\xi_{\mathbf{w}},\eta_{\mathbf{w}},\Phi_{\mathbf{w}},g_{\mathbf{w}}\right)$ which is the restriction of the weighted Sasaki structure on the sphere $S^{2n+1}$ with Reeb vector field $\xi_{\mathrm{w}}=\sum_{k=0}^{n}w_{k}\left(y_{k}\partial_{{x}_{k}}-x_{k}\partial_{{y}_{k}}\right)$ and contact form $\eta_{\mathbf{w}}=\frac{\eta}{\sum_{i=0}^{n}w_{i}\left(\left(x_{i}\right)^{2}+\left(y_{i}\right)^{2}\right)},$ where $\eta$ denotes the standard contact 1-form on the sphere $S^{2n+1}.$ If one considers the locally free $S^{1}$-action induced by the weighted $\mathbb{C}^{*}$ action on $V_{f},$ the quotient space of the link $L_{f}(\mathbf{w},d)$ by this action is the weighted hypersurface $Z_{f},$ a Kähler orbifold. We have the following commutative diagram [9]